Player Character Sheet for use with AD&D 2^{nd} Edition

Player Name:  Character Name:  Sex:
Race:  Age:  Height:  Weight:  Hair:  Eyes:
Class:  Class Kit:  Handedness:
Social Class:  Birth Rank:  Siblings: M  F

Religion: Alignment:

Level:  Experience Points:  Experience for Next Level:
Strength:  Hit:  Damage:  Weight:  Press:  Doors:  Bars/Gates:
Dexterity:  Reaction Adj:  Missile Attack Adj:  Defense (AC) Adj:
Constitution:  HP Adj:  System Shock:  Resurrection Survival:  Poison Save:  Regen:
Intelligence:  # of Lang/Prof.:  Spell Level:  Learn Spell:  Spells/Level:  Immunity:
Wisdom:  Magic Defense:  Bonus Spells:  % Fail:  Immunity:
Charisma:  Max. # Henchmen:  Loyalty Base:  Reaction Adj:

Comeliness: Description:

Racial Bonuses and Abilities:

Hit Points:  Hit Dice Type: d  Attacks/Round:  Movement Rate:
To Hit AC:  10 9 8 7 6 5 4 3 2 1 0 -1 -2 -3 -4 -5 -6 -7 -8 -9 -10

Roll gtr or eql:

Armor Type:  AC:  Other:  AC:
Shield Type:  AC:  Total Armor Class:
Helm Type:  AC:  w/o Shield:  w/o Dexterity:
Saving Throw  Base  Modifier  Total

Paralyze/Poison/Death Magic Rod/Staff/Wand Petrification/Polymorph Breath Weapon Spells

Special Abilities and Restrictions:

Languages:

Spell Progression: Specialist:

Wizard: 1^{st}:  2^{nd}:  3^{rd}:  4^{th}:  5^{th}:  6^{th}:  7^{th}:  8^{th}:  9^{th}:
Priest: 1^{st}:  2^{nd}:  3^{rd}:  4^{th}:  5^{th}:  6^{th}:  7^{th}:  8^{th}:  9^{th}:

Player Character Sheet for use with AD&D 2^{nd} Edition (pg. 2)

Class Kit Bonus Proficiencies:

Class Kit Benefits:

Class Kit Hinderances:

Weapon Proficiency  Slots

Fighting Style Specialization  Slots

Single-Weapon Two-Handed Weapon and Shield Two-Weapons Wrestling Punching Martial Arts Ambidexterity Other: Other:

Nonweapon Proficiency  Slots Ability Stat Check Modifier Total

Weapon Carried  Cost Weight Size/Type Weap. Sp. + Dex. Adj. + Magic  Damage S–M/L

Equipment  Weight @/Encum. Purpose/Effects  Where Kept

Thieving  Pick Open Find/Remove Move Hide In Detect Climb  Read
Skills  Pockets Locks  Traps  Silently Shadows Noise Walls Languages

Base Score: Class Kit: Racial Adj: Dex Adj: Armor Adj: Point Dist: TOTAL:

Level: 1 2 3 4 5 6 7 8 9 10 Hit Points: Current Max. Hit Points:

◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯
◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯

◯ ◯ ◯ ◯ ◯ ◯ ◯ ◯ ◯ ◯

Armor Worn: Max. Damage Points:

◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯
◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯  ◯ ◯ ◯ ◯ ◯

Wealth:

Magic Items:

Character Description, Notes (XP’s, Adventure Log, etc.), or Continuations: